Required information (The following information applies to the questions displayed below.) Part 2 of 4 At the beginning of YeCash Merchandise Inventory Beg. Bal 3b. 3a. Beg. Bal 3a. 1a. 6,900 X 11,858 7,000 X 1,500 X 190 4,752 1,680 140 15,000 5,200Common Stock Retained Earnings Beg. Bal Beg. Bal End. Bal End. Bal Sales Revenue Cost of Goods Sold Beg. Bal Beg. Bal Bal BalTransportation-out Interest Revenue Beg. Bal Beg. Bal Bal Bal Gain on Sale of Land Beg. Bal Bal
Answer #1

Posting the beginning balances and the events to the T-Accounts.

Cash a/c

Beg. Bal $6,900
7.b $11,858 $190 1.b
8 $8,500 $4,752 3.b
$1,680 5.a
$140 6
Ending bal $20,496 Ending bal

Merchandise inventory a/c

Beg. Bal $15,000
1.a $5,200 $400 2
1.b $190 $48 3a
5.b $900 $6,800 4.b
Bal $14,042
$642 10
Ending bal $13,400 Ending bal

Accounts receivable

Beg. Bal
4a $12,100 $242 7a
$11,858 7b
Ending bal Ending bal

Interest receivable

Beg. Bal
9 $600   
Ending bal $600 Ending bal

Land a/c

Beg. Bal $7,000
$7,000 8   
Ending bal Ending bal

Accounts payable

Beg. Bal
2 $400 $5,200 1a
3a $48
3b $4,752
Ending bal Ending bal

Common stock a/c

Beg. Bal $15,000
Ending bal $15,000 Ending bal

Retained earnings a/c

Beg. Bal $13,900
$2,100 cl.
Cl. $6,682 $10,170 cl.
Ending bal $19,488 Ending bal

Sales Revenue a/c

Beg. Bal
5a $1,680 $12,100 4a
7a $242
Ending bal $10,170 Ending bal
Cl. $10,170

Cost of goods sold a/c

Beg. Bal
4b $6,800 $900 5b
10 $642
Ending bal $6,452 Ending bal
$6,452 Cl.

Transportation out

Beg. Bal
6 $140
Ending bal $140 Ending bal
$140 Cl.

Interest revenue a/c

Beg. Bal
$600 9
Cl. $600
Ending bal Ending bal

Gain on sale of land

Beg. Bal
$1,500 8
Cl. $1,500
Ending bal Ending bal

Notes:

3a = Accounts payable (5,209-400×0.01)=$48

3b = Accounts payable (5,200-409-48)=$4,752

7a = Accounts receivable (12,100×0.02)=$242

7b =Accounts receivable (12,100-242)=$11,858

____×____

All the best

Please do support